OpenStudy (anonymous):

what is the next three likely terms in this sequence 40,4,0.4,0.04

OpenStudy (anonymous):

keep dividing by 10. 0.04/ 10 = 0.004, 0.004/10 = 0.0004, 0.0004/10 = 0.00004